Bob and Wendy live in a house in which the floor of every room is a square and covered with identical square tiles. Wendy's room contains 101 tiles more than Bob's room. How many tiles does Wendy's room contain?
Answer by AnlytcPhil(1806)

Bob and Wendy live in a house in which the floor of every 
room is a square and covered with identical square tiles. 
Wendy's room contains 101 tiles more than Bob's room. How 
many tiles does Wendy's room contain?

Let x = the number of tiles along one side of Wendy's room.
Then x² = the number of tiles in Wendy's room.

Let y = the number of tiles along one side of Bob's room.
Then y² = the number of tiles in Bob's room.

>>...Wendy's room contains 101 tiles more than Bob's 
room...<<

Translation:  x² = y² + 101

              x² - y² = 101      
               
            (x - y)(x + y) = 101

This says (x - y) and (x + y) are both factors of 101

But 101 is a prime number and thus has only two factors, 
1 and 101.

Therefore x - y, the smaller one, must equal 1, and 
x + y, the larger one, must equal 101. So we solve the 
system of equations:

x - y = 1
x + y = 101

and we get x = 51, y = 50

So Wendy's room contains x² or 51² or 2601 tiles
and Bob's room contains y² or 50² or 2500 tiles.

Edwin